Erie Company manufactures a small mp3 player called the Jogging Mate. The company uses standards to control its costs. The labor standards that have been set for one Jogging Mate mp3 player are as follows:

Standard
Hours   Standard Rate
per Hour   Standard
Cost
30 minutes   $6.00   $3.00

During August, 10,380 hours of direct labor time were needed to make 19,200 units of the Jogging Mate. The direct labor cost totaled $61,242 for the month.

Required:
1.  
According to the standards, what direct labor cost should have been incurred to make 19,200 units of the Jogging Mate? By how much does this differ from the cost that was incurred? (Round Standard labor time per unit to 2 decimal places.)
      
  
2.  
Break down the difference in cost from (1) above into a labor rate variance and a labor efficiency variance. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round intermediate calculations.)
  
3.  
The budgeted variable manufacturing overhead rate is $4.7 per direct labor-hour. During August, the company incurred $58,128 in variable manufacturing overhead cost. Compute the variable overhead rate and efficiency variances for the month. (Indicate the effect of each variance by selecting "F" for favorable, "U" for unfavorable, and "None" for no effect (i.e., zero variance). Do not round intermediate calculations and round your final answers to nearest whole dollar.)
Answer #1

1) Schedule

No of units manufactured 19200
Standard labor hour per unit 0.50
Standard labor hour allowed 9600
Standard rate per hour 6
Total Standard cost 57600
Actual labor cost 61242
Total Standard labor cost 57600
Total labor variance 3642 U

2) Labor rate variance = (6*10380-61242) = 1038 F

Labor efficiency variance = (9600-10380)*6 = 4680 U

3) Variable overhead rate variance = (4.7*10380-58128) = 9342 U

Variable overhead efficiency variance = (9600-10380)*4.7 = 3666 U